Make New Entry,
Make Followup Entry
User name N Toro
Log entry time 00:01:07 on July 6,2010
Entry number 330355
keyword=July 5 Swing shift summary
Swing shift summary:
Shift Leader: Natalia Toro
TO/Replay: Ahmed Zafar
Run 1750: 1M events 0.5uA production on Pb #1
Accepting all T6, some T1,T3,T4 since T6 rate is much smaller than bandwidth
1751: same (700K events)
17:36 Moving target to optics
1752: spot++ run on optics BPMs at (0,0)
At this position, we seem to graze optics frame.
1754,1755: tune beam, BPMs at (0,-2)
1756: tune beam, BPMs at (0,-3)
Above 4 runs have 12x8 MCC raster
1758: lowered raster to 0.8x8 in MCC at (0,-3) BPMs, tune beam.
Lowering raster to 0.8x4 MCC.
1760: bpm a,b=(0,-3)
1762: bpm A, B=(0,-5)
1764: bpm A, B=(0,-4) with RASTER OFF
1767: coincidence data on thick carbon (5uA CW): T6 rate=33 Hz.
20:14 -- taking escorted access to modify T6 as right-arm positron singles trigger, and insert sieve for optics runs.
Moved to Ta target.
run 1769: Optics data with sieve in, raster off and 5uA on single-arm modified T6.
22:03 resuming beam; recalibrating ion chambers on Ta up to 60uA.
Raster is 2x2 mcc units.
Raising raster to 12x8
Production run: 1778 at 60 uA
1779: test run for spot++ at 1uA with 6x6 MCC raster
1780: production run at 6x6 MCC raster, 60 uA, having trouble analyzing events.
1781: coda not functioning
1783: production at 6x6 MCC, 60uA.
Spot++ confirms we are no longer hitting any frame. The coincidence rate is low (18 Hz) but it is almost all true coincidences.
23:55 Q1 down; we are requesting controlled access. Pete is going into the hall to clear the fault.